# AI for Just Work: Constructing Diverse Imaginations of AI beyond “Replacing Humans”

### v1

@@ 1. The phenomenon of monoculture in the AI community, and why questioning the objectives of AI matters?
Regarding the vision that “AI can fix its own flaws,” we refer to the following quote and critique from Jenna Burrell’s paper Automated decision-making as domination (Burrell, 2024):
‘In early 2020, just before the COVID-19 pandemic shut down all large gatherings for the forseeable future, the Association for the Advancement of Artificial Intelligence hosted its annual event. At a keynote panel, AI researchers Yoshua Bengio, Geoffrey Hinton, and Yann LeCun, as recent winners of the Turing Award for lifetime achievement in computing, were given the stage to speak about their technical breakthroughs. A headline summarized the gist of their talks this way: “Deep learning godfathers Bengio, Hinton, and LeCun say the field can fix its flaws” (Ray, 2020). This sentiment captures the essence of disciplinary conservativism. The ritual of posing questions about the future of a research field to its most senior and successful members is not peculiar to AI research or computer science as a discipline. These are realities of how any field that has garnered resources protects its claims to legitimacy (Fourcade, et al., 2015). Yet, one can also imagine the alternatives. What if these senior scholars were to say, “… the field needs outside regulators to set limits” or “… the field requires independent auditors to cast a fresh eye on its flaws” or “Deep learning godfathers … cede the floor to promising early career scholars ….” While the ostensible purpose of academic fields is to generate new knowledge, they are at the same time social systems; designed to protect resources and aimed ultimately at social reproduction and self-perpetuation.’ [^20] From Burrell (2024) Section 1. Introduction: “The field can fix its flaws”
This vision often reflects the techno-solutionism or technological fix mindset that ignores the complexity of the problem space and resort to technology as the only source of solution  (Morozov, 2014; Tec, 2024).

### v2

@@ 1. The phenomenon of monoculture in the AI community, and why questioning the objectives of AI matters?
Regarding the vision that “AI can fix its own flaws," we refer to the following quote and critique from Jenna Burrell’s paper Automated decision-making as domination [30]:
‘In early 2020, just before the COVID-19 pandemic shut down all large gatherings for the forseeable future, the Association for the Advancement of Artificial Intelligence hosted its annual event. At a keynote panel, AI researchers Yoshua Bengio, Geoffrey Hinton, and Yann LeCun, as recent winners of the Turing Award for lifetime achievement in computing, were given the stage to speak about their technical breakthroughs. A headline summarized the gist of their talks this way: “Deep learning godfathers Bengio, Hinton, and LeCun say the field can fix its flaws” (Ray, 2020). This sentiment captures the essence of disciplinary conservativism. The ritual of posing questions about the future of a research field to its most senior and successful members is not peculiar to AI research or computer science as a discipline. These are realities of how any field that has garnered resources protects its claims to legitimacy (Fourcade, et al., 2015). Yet, one can also imagine the alternatives. What if these senior scholars were to say, “… the field needs outside regulators to set limits” or “… the field requires independent auditors to cast a fresh eye on its flaws” or “Deep learning godfathers … cede the floor to promising early career scholars ….” While the ostensible purpose of academic fields is to generate new knowledge, they are at the same time social systems; designed to protect resources and aimed ultimately at social reproduction and self-perpetuation.’ [^22] From Burrell, [30] Section 1. Introduction: “The field can fix its flaws”
This vision often reflects the techno-solutionism or technological fix mindset that ignores the complexity of the problem space and resort to technology as the only source of solution  [79, 108].